Solve
a)log(base 2)(x + 1)+ log(base 2)(x-1) = 3
b)log(base 2)(x^2-1) = 3
c)explain why the equations in part a and b are not equivalent
Can you please help me out? Thanks so much in advance:)

Answer by edjones(8007)

a)
log(base 2)(x + 1)+ log(base 2)(x-1) = 3
2^log(base 2)(x + 1)*2^log(base 2)(x-1)=2^3
(x+1)(x-1)=8
x^2=9
x=3
.
b)
log(base 2)(x^2-1) = 3
2^log(base 2)(x^2-1) =2^3
x^2-1=8
x^2=9
x=3, x=-3
.
c) x can't equal -1 in a) because a log cannot equal 0 or be negative.
.
Ed